OpenStudy (anonymous):

what is the solution of the equation? (x+6) to the 3/5 =8

OpenStudy (anonymous):

\[(x+6)^{3/5}=8\]

OpenStudy (anonymous):

x+6

OpenStudy (anonymous):

answer: 14, 2 ,26 ,38

OpenStudy (anonymous):

(x+6)^(3/5) = 8 (x+6)^(3) = 8^5 x+6 = 8^(5/3) x = 8^(5/3) - 6

OpenStudy (anonymous):

Let me know if you need further explanation @shanay_94

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

What's confusing?

OpenStudy (anonymous):

like i dont get the raising to the power thing

OpenStudy (anonymous):

\[(x^a)^b = x^{a*b}\]

OpenStudy (anonymous):

so we can take (x+6)^(3/5) = 8 and put that to the fifth power on both sides: ((x+6)^(3/5))^5 = 8^5 using the identity i stated, this can simplify to (x+6)^((3/5) * 5) = 8^5 which simplifies to (x+6)^3 = 8^5

OpenStudy (anonymous):

then you can use the same process by putting both sides to the 1/3 power (cube root) to get x+6 = (8^5)^(1/3) x+6 = 8^(5/3)